a bullet is fired horizontally above a tunnel 6m high and strikes a target 100m away. Find the time if flight. (g=10ms-2)
1
Expert's answer
2019-06-10T09:59:25-0400

We have


"h=0.5gt^2"

The time of flight:


"t=\sqrt{\frac{2h}{g}}"

"t=\sqrt{\frac{2(6)}{10}}=1.1\ s"